OpenStudy (anonymous):

Let A,B⊆R be nonempty. Define A−B={a−b|a∈A,b∈B}. Prove that inf(A−B)=inf(A)−sup(B)

OpenStudy (anonymous):

Any ideas @Zarkon or @dirtydan667 ?

OpenStudy (zarkon):

yes...I know how to prove it...I wanted to see if you have anything.

OpenStudy (anonymous):

no im drawing a blank

OpenStudy (anonymous):

I have nothing

OpenStudy (zarkon):

can you prove...inf(A-B)=inf(A)+inf(-B)?

OpenStudy (zarkon):

then show that inf(-B)=-sup(B)

OpenStudy (zarkon):

both are standard results

OpenStudy (anonymous):

I can prove inf(A+B)=inf(A)+inf(B)

OpenStudy (anonymous):

Is the subtraction the same proof or no?

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

You still there @Zarkon ?

OpenStudy (anonymous):

no i killed him

OpenStudy (anonymous):

@Zarkon can you please help me finish the proof?

OpenStudy (zarkon):

if you can prove inf(A+B)=inf(A)+inf(B) then That is the same as inf(A-B)=inf(A)+inf(-B) let C=-B then inf(A-B)=inf(A)+inf(-B) is the same as inf(A+C)=inf(A)+inf(C)

OpenStudy (anonymous):

And then I show that inf(-b) is same as sup(b)

OpenStudy (anonymous):

?